Session: Oct 2005 (session 3)
Wavelength: 21 cm


To give feedback on this experiment, click here

Experiment Feedback

Torun (Kaz Borkowski at Fri Nov 4 13:11:33 2005) Failure :
Not observed due to lack of disk capacity. One of the 8-packs provided by JIVE proved defective and new disks purchased by us (to fill an empty module) arrived last minute but most unfortunately were found to be of a wrong type. (The remaining US destined GM057B and 0.16 TB EA035A will be recorded as scheduled.)

Medicina (Andrea Orlati at Fri Nov 4 21:27:22 2005) Success :

Onsala (Michael Lindqvist at Sat Nov 5 07:24:57 2005) Minor failures :
Very strong RFI affected (overflow) two BBCs (5&6) from around 18:10 UT. Otherwise no known problems.

Westerbork (Tony Foley at Tue Nov 8 08:19:57 2005) Minor failures :
Wb used in phased array mode. RT9 removed from the array as there were delay hardware problems for this telescope in one band. Hence 13 telescopes.

Effelsberg (A: Kraus at Wed Nov 9 16:41:12 2005) Success :

Jodrell_Bank (Alastair Gunn at Thu Nov 10 12:37:00 2005) Minor failures : ANT
The Lovell Telescope was used. A serious azimuth drive failure occurred at about 13:11 UT. Was repaired and back on source at about 18:19 UT. No other known problems.